# Steering Large Language Models for Machine Translation Personalization

### v1

@@ Abstract
High-quality machine translation systems based on large language models (LLMs) have simplified the production of personalized translations reflecting specific stylistic constraints. However, these systems still struggle in settings where stylistic requirements are less explicit and might be harder to convey via prompting. We explore various strategies for personalizing LLM-generated translations in low-resource settings, focusing on the challenging literary translation domain. We explore prompting strategies and inference-time interventions for steering model generations towards a personalized style, and propose a contrastive framework exploiting latent concepts extracted from sparse autoencoders to identify salient personalization properties. Our results show that steering achieves strong personalization while preserving translation quality. We further examine the impact of steering on LLM representations, finding model layers with a relevant impact for personalization are impacted similarly by multi-shot prompting and our steering method, suggesting similar mechanism at play. [^1] Code available at DanielSc4/steering-for-personalization.
{CJK*}
UTF8gbsn
@@ 1 Introduction
Figure 1: We compare prompt-based approaches with steering techniques intervening on model internals for personalizing MT outputs in literary machine translation, employing MT quality metrics and style classifiers to disentangle the effect of steering on outputs fluency and personalization adequacy.
When we read a translated book, we do not simply read the story in a new language; we also experience the translator’s personal voice by means of their stylistic choices. Past efforts in the automatic translation of literary works have historically been constrained by the limited capabilities and flexibility of machine translation (MT) systems. The recent popularization of MT systems based on large language models (LLMs) has greatly improved their capacity in handling the long contexts typical of literary translations, but mimicking the creative and rich language that characterize the translators’ own style remains an open issue. In this context, several works explored the usage of prompting and tuning-based strategies to ensure translations are stylistically appropriate (Michel and Neubig, 2018; Wang et al., 2021). However, their influence on model internal representations is rarely explored, so that their impact is less controllable and often unpredictable. In this work, we adopt prompting methods as baselines and compare their effectiveness to steering methods proposed in interpretability literature. These techniques can be used to surgically intervene on LLMs’ intermediate representation to generate personalized translations when few examples are available. For this purpose, we employ the par3 dataset (Thai et al., 2022), which includes multiple human translations for novels translated to English from 7 typologically diverse languages.
@@
We begin with preliminary assessments by verifying whether translators’ styles are discernible by automatic systems, finding that trained classifiers can distinguish writing styles with high accuracy, while the task is notoriously challenging for human annotators (Youyou et al., 2015; Flekova et al., 2016). We also find a simple prompting setting with in-context personalization examples to improve the style accuracy of LLM translation, suggesting personalized translation styles are reproducible. We connect the conditioning induced by prompting to the inner workings of the model, identifying activations with high discriminative capacity for style differences in intermediate model layers. We then propose a contrastive steering approach based on sparse autoencoders (SAEs, Cunningham et al., 2023) to condition model generations by upweighting sparse, interpretable latents at inference time. We validate the effectiveness of our method across three LLMs of various sizes on par3 novels, comparing our results with established prompting and steering methods.
@@
Our results show that contrastive SAE steering is a promising approach for MT personalization, leading to translations that are not only more in line with general human translation features but also more aligned with the desired personalized style compared to other methods. Importantly, these results are achieved with no translation quality degradation according to established MT quality metrics. We conclude by comparing the impact of our method on model representations with the outcome of multi-shot prompting, finding that probes trained on prompt-conditioned activations can predict the effectiveness of SAE steering with high precision. These results confirm that tested prompting and steering techniques converge to similar solutions for conditioning model behavior, enabling future investigations into the mechanistic impact of prompting through the study of learned SAE latents and other interpretable components.
@@ 2 Related Work
The literary domain has historically been challenging for automatic MT systems due to their limited ability in handling rich linguistic and cultural contexts (Matusov, 2019) and their propensity to produce overly literal outputs (Guerberof-Arenas and Toral, 2022). Automatic literary translation has a long history dating back to pre-neural MT approaches (Voigt and Jurafsky, 2012; Toral and Way, 2015, 2018; Moorkens et al., 2018) with two recent dedicated evaluation campaigns (Wang et al., 2023b, 2024a). The advent of LLMs brought new opportunities in the processing of longer context for document-level translation (Wang et al., 2023a; Briakou et al., 2024; Wu et al., 2025a), but critical errors requiring human translator’s intervention nonetheless persist (Karpinska and Iyyer, 2023). In this work, we use the par3 dataset (Thai et al., 2022) containing multiple human translations of novels to evaluate MT personalization in the literary domain.
@@
Advances in MT quality recently led to a growing interest in personalization approaches to ensure a consistent format and appropriate stylistic choices in model generations (Rabinovich et al., 2017; Lin et al., 2021). Previous approaches for controlling attributes such as formality (Sennrich et al., 2016; Niu et al., 2017; Nadejde et al., 2022) or gender (Vanmassenhove et al., 2018; Saunders and Byrne, 2020) typically required tuning existing models on pre-defined properties of interest, with few works attempting a real data-driven adaptation from unlabeled demonstrations (Michel and Neubig, 2018; Wang et al., 2021; Zhang et al., 2022). More recently, several studies employed prompting (Garcia and Firat, 2022; Sarti et al., 2023) or preference optimization from post-editing behavior (Lee et al., 2023; Berger et al., 2024) to render MT personalization more effective and data-efficient. In this work, we complement prompt results with steering approaches to personalize MT outputs using few user-provided examples.
@@
Steering approaches exploit the linear structure of LM activations (Mikolov et al., 2013; Chanin et al., 2024) to craft inference-time interventions for influencing model generations. These methods commonly employ contrastive sets of in-context demonstrations (Rimsky et al., 2024; Scalena et al., 2024) to map input properties to components such as vectors (Turner et al., 2024; Li et al., 2023), linear probes (Zou et al., 2025) or learned projections (Wu et al., 2024, 2025b). Sparse Autoencoders (SAEs) are another family of promising approaches for enabling fine-grained interventions in language models (Yun et al., 2021; Cunningham et al., 2023; Templeton et al., 2024). They are trained to decompose activations into approximately monosemantic features, offering a potentially interpretable basis for modifying model behavior. While interpreting their learned latents remains non-trivial (Marks et al., 2025), SAEs have proven effective for applying targeted interventions along specific linear directions (Chalnev et al., 2024; Zhao et al., 2025; Ferrando et al., 2025). However, most research on SAEs have so far focused on synthetic tasks or standard benchmarks, leaving their potential in real-world settings relatively underexplored.
@@ 3 Preliminaries
H1 "Such being the case," the Taoist acquiesced, "I am ready to follow you, whenever you please to go."
H2 "Very good, I will go with you then," said the Taoist.
Gemma 2 2B "If that’s the case, then go ahead" said the Taoist.
Gemma 2 9B "If so, then I will go with you" the Taoist said
Llama 3.1 8B The Taoist said: "If you insist on going, then go ahead."
@@
Gemma 2 2B "It will be the effect of the sea water. The sea makes of these jokes."
Gemma 2 9B It will be the effect of the sea water. The sea plays these tricks.
Llama 3.1 8B It will be the effect of the sea water. The sea does things like this.
Table 1: zh $\rightarrow$ en and it $\rightarrow$ en examples for par3 segments translated by humans (H1, H2) and LLMs with zero-shot prompting. More examples in Appendix D.
Before testing the effectiveness of personalization strategies, we validate some key assumptions: i) Whether the personalized translation style is discernible, i.e., if it is possible to tell apart human- and machine-generated translations; ii) Whether different translation styles are automatically reproducible, i.e., if LLMs can mimic a specific translator’s style when provided with some examples; and iii) Whether style distinctions are reflected in the model’s internal representations, to motivate the interest in steering approaches for personalization.
@@
3.1 Are Personalized Translations Discernible?
Following prior work on personalization (Wang et al., 2024c; Liu et al., 2023), we train a series of [Uncaptioned image] classifiers based on multilingual XLM Transformer encoders (Conneau et al., 2020) to distinguish between H1, H2, and MT translations. If those systems can reliably separate these three classes, it suggests the presence of reasonably distinct stylistic signals differentiating them. In particular, the ability to distinguish between H1 and H2 would denote not only the possibility to discern a human-like style from human-made and automatic translations, but also a personalized style from different human translators.
[…]
@@ 3.1 Are Personalized Translations Discernible?
3.2 Can LLMs Reproduce Human Translation Styles?
To confirm whether MT personalization can be achievable, we test LLMs ability to mimic the stylistic choices of a particular translator in a multi-shot (MS) prompting setup. For each translator available across tested novels, we provide the model with 20 in-context examples selected from the original pool of translated paragraphs by that translator, asking it to generate a consistent translation. We compare MS results with the default zero-shot (ZS) prompting without any example from the translator to quantify the effect of in-context examples. Table 2 presents results for personalization accuracy, automatically evaluated using our high-scoring [Uncaptioned image] classifiers from the previous section; and translation quality, estimated via the widely used [Uncaptioned image] Comet MT metric (Rei et al., 2020). The proportion of outputs categorized as matching the translator’s style is increased two- to four-fold following MS prompting, suggesting that LLMs can employ implicit clues in small sets of user examples to produce personalized translations. Stable scores for [Uncaptioned image] Comet also confirm that translation quality is maintained during style adaptations.
3.3 Finding Personalization Information in LLM Representations
@@
Figure 2: Probing classifier performance on the human translation detection task across Gemma 2 2B layers. Activations in intermediate layers are found to capture translation style information with high precision, with layer 13 performing best across all tested languages.
In light of these results, we set out to test where the model encodes information reflecting a stylistic shift when style-appropriate examples are provided. To this purpose, we train linear probes (Belinkov, 2022) using model activations as input features to predict the style label (MT, H1, or H2) that the style [Uncaptioned image] classifier (from Section 3.1) would assign to the eventual translation, based purely on the prompt’s internal representation. Probing accuracy is measured by testing their accuracy in predicting the classified outcome before the generation, using only the prompt representation formed by the model. Given a test set of human-translated paragraphs, we train our probes on a set of examples using an MS prompt with 20 in-context examples. The set is balanced between prompts showcasing personalization with gold in-context examples from a human translator, and non-personalized prompts with MT-generated examples previously produced by the same tested model in a ZS setup. Test examples are selected from the respective novels to ensure for the [Uncaptioned image] classifier prediction shifts from MT in the ZS setting to the style of in-context examples when MS is used, signaling a causal influence of demonstrations on output personalization. [^5] Examples are resampled for every test paragraph to prevent the probe from overfitting on spurious prompt features. This balanced setup prevents the leaking of task information, e.g., number of in-context examples, to learned probes, ensuring that stylistic differences among human and MT-generated in-context examples are the sole factor determining differences in model activations. We focus specifically on Gemma models, extracting activations after the attention block at each model layer for the last token of the prompt, which was previously shown to encode key task-relevant information (Hendel et al., 2023; Todd et al., 2024; Scalena et al., 2024)Figure 2 reports probe accuracies across all Gemma 2 2B layers, with results for the 9B model reported in Appendix C. We find a peak in probe accuracy of $\sim$95around intermediate model layers, suggesting that these layers encode stylistic information with near-perfect precision. [^6] We find probes for layers 13 and 21 to perform best for the 2B and 9B models, respectively. These results confirm that personalization is discernible from LLMs’ internal representation, motivating our experiments towards the design of inference-time interventions to steer models towards personalized MT outputs.
@@ 4 Methods
We begin by introducing our the prompting and steering methods that we use as baselines and outline our own proposed SAE-based steering approach for personalized translation.
@@ 4.1 Prompting Baselines
The ZS setup used in our main experiment correspond to the one from Section 3.1, in which the model is simply asked to produce a translation with no conditioning from examples or explanations towards the target translation style. We use this setting to establish a baseline style and translation quality performance for the models.
Zero-Shot Explain (ZS-Expht and ZS-Exppt).
Building upon the ZS setting, we experiment with a prompting strategy where LLMs are provided with detailed explanations (Exp) of the most salient elements that characterize the desired translation style. We obtain such descriptions by prompting a capable proprietary model, GPT-4o OpenAI (2024), with 20 translations matching the desired style, asking it to synthesize a set of guidelines to matching the examples. We evaluate two contrastive variants of this approach, providing GPT-4o with either MT examples (ZS-Expht) or alternative human translations (ZS-Exppt) alongside examples matching the desired style, and asking to describe what characterizes the latter compared to the former. To avoid data leakage, all generated explanations are manually reviewed to ensure they do not contain any verbatim content or direct excerpts from the input examples. [^7] Details on the prompt templates are in Appendix B.5. Tested models are then prompted with GPT-4o explanations in a ZS setting, to verify whether interpretable directives synthesized from a set of examples matching the desired behavior can produce reliable personalization results.
@@
Following Section 3.2’s findings, we adopt the same MS setup using 20 in-context translation examples matching the style of a target human translator (H1 or H2).

### v2

@@ Abstract
Large language models have simplified the production of personalized translations reflecting predefined stylistic constraints. However, these systems still struggle when stylistic requirements are implicitly represented by a set of examples, such as texts produced by a specific human translator. In this work, we explore various strategies for personalizing automatically generated translations when few examples are available, with a focus on the challenging domain of literary translation. We begin by determining the feasibility of the task and how style information is encoded within model representations. Then, we evaluate various prompting strategies and inference-time interventions for steering model generations towards a personalized style, with a particular focus on contrastive steering with sparse autoencoder (SAE) latents to identify salient personalization properties. We demonstrate that contrastive SAE steering yields robust style conditioning and translation quality, resulting in higher inference-time computational efficiency than prompting approaches. We further examine the impact of steering on model activations, finding that layers encoding personalization properties are impacted similarly by prompting and SAE steering, suggesting a similar mechanism at play.
@@ 1 Introduction
Figure 1: We compare prompt-based approaches with steering techniques intervening on model internals for personalizing MT outputs in literary machine translation. We use MT quality metrics and style classifiers to quantify the impact of steering on output fluency and personalization accuracy.
A translated novel is not a simple literal retelling of the story—the personal style of the translator and their lexical and stylistic choices play a crucial role in shaping the narrative in a new language. Past efforts in the automatic translation of literary works have historically been constrained by the limited capabilities and flexibility of machine translation (MT) systems. The recent popularization of MT systems based on large language models (LLMs) has dramatically improved their capacity in handling the long contexts typical of literary translations, but mimicking the creative and rich language that characterizes the translators’ own style remains an open issue. While prompting and tuning-based strategies have been successfully used for MT personalization (Michel and Neubig, 2018; Wang et al., 2021), recent advances in interpretability research have highlighted the effectiveness of various steering methods, which employ surgical interventions on LLMs’ activations to condition model generation (Rimsky et al., 2024). While steering techniques were primarily applied to explicit properties such as sentiment and formality, this work aims to test their effectiveness for MT personalization. We compare steering methods with established prompting techniques, with a particular focus on the challenging domain of literary translation, where stylistic choices are more evident.
@@
Provided that stylistic attribution is notoriously challenging for human annotators (Youyou et al., 2015; Flekova et al., 2016), we begin with a preliminary assessment to verify whether trained classifiers can distinguish between translators’ styles and whether LLMs can use demonstrations to enhance the stylistic accuracy of their translations. We then connect the conditioning induced by prompting to the inner workings of the model, identifying activations with high discriminative capacity for style differences in intermediate layers of LLMs. In light of our results indicating that style information is effectively detected and represented by automatic systems, we evaluate the effectiveness of various established prompting and steering methods on three multilingual LLMs across novels in seven mid- to high-resourced languages. We focus in particular on a contrastive steering approach using sparse autoencoders (SAEs, Cunningham et al., 2023) to condition model generations by upweighting sparse style-related latents at inference time.
@@
Our results demonstrate that contrastive SAE steering is a promising approach for MT personalization, yielding translations that align more closely with general human translation features and exhibit a greater similarity with the desired personalized style compared to other methods. Importantly, these results are achieved with no degradation in translation quality, assessed according to established MT quality metrics. We conclude by comparing the impact of our method on model representations with the outcome of multi-shot prompting, finding that probes trained on prompt-conditioned activations can predict the effectiveness of SAE steering with high precision. These results confirm that prompting and SAE steering techniques converge to similar solutions for conditioning model behavior, enabling future investigations into the mechanistic impact of prompting through the study of interpretable SAE latents.
@@ 2 Related Work
The literary domain has historically been challenging for automatic MT systems due to their limited ability in handling rich linguistic and cultural contexts (Matusov, 2019) and their propensity to produce overly literal outputs (Guerberof-Arenas and Toral, 2022). Automatic literary translation has a long history dating back to pre-neural MT approaches (Voigt and Jurafsky, 2012; Toral and Way, 2015, 2018; Moorkens et al., 2018) with two recent dedicated evaluation campaigns (Wang et al., 2023b, 2024a). The advent of LLMs brought new opportunities in the processing of longer context for document-level translation (Wang et al., 2023a; Briakou et al., 2024; Wu et al., 2025a), but critical errors requiring human translator’s intervention nonetheless persist (Karpinska and Iyyer, 2023). In this work, we use the par3 dataset (Thai et al., 2022), which comprises multiple human translations of novels to evaluate MT personalization in the literary domain.
@@
Advances in MT quality recently led to a growing interest in personalization approaches to ensure a consistent format and appropriate stylistic choices in model generations (Rabinovich et al., 2017; Lin et al., 2021). Previous techniques for controlling attributes such as formality (Sennrich et al., 2016; Niu et al., 2017; Nadejde et al., 2022) or gender (Vanmassenhove et al., 2018; Saunders and Byrne, 2020) typically required tuning existing models on pre-defined properties of interest, with few works attempting a real data-driven adaptation from unlabeled demonstrations (Michel and Neubig, 2018; Wang et al., 2021; Zhang et al., 2022). More recently, several studies employed prompting (Garcia and Firat, 2022; Sarti et al., 2023) or preference optimization from post-editing behavior (Lee et al., 2023; Berger et al., 2024) to render MT personalization more effective and data-efficient. In this work, we complement prompt results with steering approaches to personalize MT outputs using a few user-provided examples.
@@
Steering approaches exploit the linear structure of LM activations (Mikolov et al., 2013; Chanin et al., 2024) to craft inference-time interventions for influencing model generations. These methods commonly employ contrastive sets of in-context demonstrations (Rimsky et al., 2024; Scalena et al., 2024) to map input properties to components such as vectors (Turner et al., 2024; Li et al., 2023), linear probes (Zou et al., 2025), or learned projections (Wu et al., 2024, 2025b). Sparse Autoencoders (SAEs) are another family of promising approaches for enabling fine-grained interventions in language models (Yun et al., 2021; Cunningham et al., 2023; Templeton et al., 2024). They are trained to decompose activations into approximately monosemantic features, offering a potentially interpretable basis for modifying model behavior. While interpreting their learned latents remains non-trivial (Marks et al., 2025), SAEs have proven effective for applying targeted interventions along specific linear directions (Chalnev et al., 2024; Zhao et al., 2025; Ferrando et al., 2025). However, most research on SAEs has so far focused on synthetic tasks or standard benchmarks, leaving their potential in real-world settings relatively underexplored.
@@ 3 Preliminaries
Lang. Novel Train / Val / Test
IT Pinocchio 745 / 82 / 107
FR Around the World in Eighty Days 829 / 92 / 120
NL The Diary of a Young Girl 769 / 85 / 110
DE Beware of Pity 606 / 67 / 96
RU Crime and Punishment 1517 / 168 / 224
JA No Longer Human 652 / 40 / 81
ZH Dream of the Red Chamber 694 / 76 / 92
Table 1: Novels and number of example paragraphs used in our evaluation. Each example corresponds to a triple (MT, H1, H2). 20 additional examples per language are reserved for prompting/steering purposes.
@@
H1 “Such being the case,” the Taoist acquiesced, “I am ready to follow you, whenever you please to go.”
H2 “Very good, I will go with you then,” said the Taoist.
Gemma 2 $_{\text{2B}}$ “If that’s the case, then go ahead” said the Taoist.
Gemma 2 $_{\text{9B}}$ “If so, then I will go with you” the Taoist said
Llama 3.1 $_{\text{8B}}$ The Taoist said:“If you insist on going, go ahead.”
@@
Gemma 2 $_{\text{2B}}$ "It will be the effect of the sea water. The sea makes of these jokes."
Gemma 2 $_{\text{9B}}$ It will be the effect of the sea water. The sea plays these tricks.
Llama 3.1 $_{\text{8B}}$ It will be the effect of the sea water. The sea does things like this.
Table 2: Examples from par3 for zh $\rightarrow$ en (“Dream of the Red Chamber” by Cao Xueqin) and it $\rightarrow$ en (“The Adventures of Pinocchio” by Carlo Collodi), including two human translations (H1, H2) and LLM outputs with zero-shot prompting. More examples in Appendix˜C.
We begin our investigation by validating some key assumptions: i) Whether the personal translation styles are discernible, i.e., if the translation style of various LLMs and human translators can be consistently identified; ii) Whether LLMs can mimic specific translators’ styles when provided with some of their translations; and iii) Whether style distinctions are reflected in the model’s internal representations, to motivate the interest in steering approaches for improving MT personalization.
@@
3.1 Is MT Personalization Discernible?
Following prior work on personalization (Wang et al., 2024c; Liu et al., 2023), we train a series of [Uncaptioned image] classifiers based on multilingual XLM Transformer encoders (Conneau et al., 2020) to distinguish between H1, H2, and MT translations for each language and each MT model in our evaluation suite. If classifiers can accurately predict the translation style among these three styles, then this provides evidence for distinct stylistic signals between them. In particular, the ability to distinguish between H1 and H2 would suggest that style signals can help differentiate between different human translators, rather than simply between human-like and automatic translations. [^1] Details in Section A.2. Full results in Section B.1. We find that the style information in the provided examples can be readily identified, with all [Uncaptioned image] classifiers reaching an accuracy between 77% for Japanese and 99% for Chinese, or 86% on average. These results corroborate previous evidence on the ability of automatic systems in detecting stylistic clues, while the task remains elusive for human annotators (Youyou et al., 2015; Flekova et al., 2016; Wang et al., 2024b).
3.2 Can LLMs Mimic Translation Styles?
We test the ability of LLMs to mimic the stylistic choices of a particular translator in a traditional multi-shot (MS) prompting setup. For each translator available across all tested novels, we prompt the model with 20 in-context examples selected from the original pool of translated paragraphs by that translator, asking it to generate new translations matching the provided style. We compare MS results with the default zero-shot (ZS) prompting using our high-scoring [Uncaptioned image] classifiers to quantify the effect of in-context examples on personalization accuracy, and using the popular [Uncaptioned image]  Comet MT metric (Rei et al., 2020) to track translation quality. Table˜3 presents our results. The proportion of outputs categorized as matching the translator’s style is increased two- to four-fold when in-context examples are provided, confirming that LLMs can employ implicit clues in small sets of user examples to improve the style accuracy of their translations. Stable scores for [Uncaptioned image]  Comet also indicate that translation quality is preserved after style adaptation.
@@ 3.2 Can LLMs Mimic Translation Styles?
3.3 Is Style Detectable in Model Activations?
@@
Figure 2: Probing classifier performance on the human translation detection task across Gemma 2 2B (top) and 9B (bottom) layers. Activations in intermediate layers are found to capture translation style information with high precision.
In light of these results, we set out to test where the stylistic information enabling the improved performance of MS prompting is encoded within model representations. To this purpose, we train linear probes (Belinkov, 2022) using the last prompt token activations as input features to predict a binary class (MT or Human) and comparing it to the label the [Uncaptioned image] classifier (from Section˜3.1) would assign to the resulting LLM translation. [^2] To simplify the probing setup, the probe Human class is used for both H1 and H2 classifier predictions. High probing accuracy in this setting would suggest that the style information encoded from prompt examples is sufficient to predict the appropriateness of model translations before they are even generated. The training set of our probes comprises examples with 20 in-context demonstrations, where half of the examples use personalized translations made by a human translator, and the other half uses MT-generated demonstrations we pre-generated using the same model in a ZS setup. For each novel, we select as test examples only left-out entries for which the [Uncaptioned image] classifier predictions flip from MT in ZS setting to the desired style (H1or H2) when demonstrations are provided, to focus our analysis on settings where in-context demonstrations produce a tangible effect on output personalization. [^3] In-context demonstrations are resampled for every test paragraph to prevent overfitting on spurious features. This balanced setup precludes shortcut learning from spurious information to trained probes, e.g., the number of prompt demonstrations, ensuring that stylistic differences between human-made and MT-generated demonstrations are the sole distinguishing factor leading to discrepancies in model activations.
We focus specifically on Gemma models, extracting activations after the attention block at each model layer for the last token of the prompt, which was previously shown to encode key task-relevant information (Hendel et al., 2023; Todd et al., 2024; Scalena et al., 2024)Figure˜2 reports probe accuracies across all Gemma 2 2B layers, with results for the 9B model reported in Appendix˜B. We find a peak in probe accuracy of $\sim$95% around intermediate model layers, suggesting that these layers encode stylistic information with very high precision. [^4] We find probes for layers 13 and 21 to perform best for the 2B and 9B models, respectively. These results confirm that personalization features are clearly discernible from LLMs’ activations, motivating our following experiments on activation steering for MT personalization.
@@ 4 Methods
We now introduce the methods evaluated in our experiments, with a focus on our proposed SAE-based contrastive steering method. In all our contrastive formulations, we treat the available H1 and H2 translations as distinct personalization targets to contrast with MT, and use $H^{+},H^{-}$ to indicate desired and undesired human styles, respectively. [^5] E.g., when mimicking H1 style, $H^{+}$H1 and $H^{-}$H2.
@@ 4.1 Prompting Baselines
The ZS setup of our main experiment matches the one from Section˜3.1, in which the model is asked to produce a translation with no conditioning from examples or explanations towards the target translation style. We use this setting to establish a baseline style and translation quality performance for tested models.
Zero-Shot with Explanation (ZS-Exp).
We extend the ZS setting by exploiting verbalized style description produced by GPT-4o OpenAI (2024). We prompt GPT-4o with 20 $H^{+}$ demonstrations and 20 MT outputs produced in the ZS settings, asking it to synthesize a detailed explanation of salient elements characterizing the desired style. [^6] Prompt templates in Listings LABEL:lst:explain-template, LABEL:lst:explain-example,LABEL:lst:explain-final of Section A.5. We manually review all generated explanations to ensure they do not contain any verbatim excerpts from input examples, and use them to the tested LLM in a setting with no demonstrations (ZS). This setup allows us to verify whether natural language directives can act as effective and interpretable alternatives to other tested methods.
@@
Following Section˜3.2’s findings, we adopt the same MS setup using 20 in-context translation examples matching the style of a desired human translator $H^{+}$.